# Model Merging in the Era of Large Language Models: Methods, Applications, and Future Directions

### 10 Mar 2026 v1

@@ 3.4 Geometric Interpolation and Manifold-Aware Methods
Since Euclidean assumptions become problematic for divergent models, geometric interpolation techniques offer principled alternatives that account for the curved structure of parameter spaces. When combining two parameter vectors $\theta_{A}$ and $\theta_{B}$ via linear interpolation $\theta_{\text{merged}}=(1-\alpha)\theta_{A}+\alpha\theta_{B}$, the resulting vector’s magnitude typically satisfies $\|\theta_{\text{merged}}\|<(1-\alpha)\|\theta_{A}\|+\alpha\|\theta_{B}\|$ unless the vectors are perfectly aligned, leading to systematic magnitude shrinkage that can degrade model representations. This phenomenon proves particularly acute when merging models with substantially different weight configurations.
Spherical Linear Interpolation (SLERP) addresses magnitude preservation by constraining the interpolation path to lie on the hypersphere connecting two points in weight space. For parameter vectors normalized to unit length, SLERP computes the interpolated point as:
@@
where $\Omega=\arccos(\theta_{A}\cdot\theta_{B}/\|\theta_{A}\|\|\theta_{B}\|)$ represents the angular distance between the two vectors. This formulation ensures that the interpolated vector maintains constant magnitude throughout the interpolation trajectory, traversing a geodesic arc rather than cutting through the interior of the sphere. Empirical investigations have demonstrated that SLERP-based merging preserves representational geometry more faithfully than linear averaging, particularly for attention weight matrices where directional information carries semantic significance.
@@
In practice, applying SLERP to neural network merging requires careful consideration of granularity. Layer-wise application, where SLERP operates independently on each layer’s parameter tensor, has emerged as the dominant approach, as global application across flattened parameter vectors fails to capture the hierarchical structure of deep networks. observed that layer-wise (Yang et al., 2024c) SLERP achieves superior capability retention compared to uniform linear averaging when merging models fine-tuned on semantically distant tasks, attributing this advantage to the preservation of layer-specific activation statistics.
@@
Beyond spherical interpolation, manifold-aware methods encompass techniques that explicitly model the non-Euclidean structure of parameter spaces. Latent space interpolation approaches, as explored in (Li et al., 2023a), project weight configurations into lower-dimensional manifolds where interpolation can be performed with reduced interference before mapping back to the original parameter space. This paradigm acknowledges that the effective dimensionality of viable weight configurations is substantially lower than the ambient parameter count, enabling interpolation paths that remain within the manifold of high-performing solutions.
@@
The variance preservation problem presents additional challenges beyond magnitude shrinkage. When averaging independent parameter vectors, the variance of the resulting distribution decreases proportionally to the number of models merged, potentially collapsing learned features toward uninformative mean values. Rescaling strategies that adjust merged parameters to match the original variance statistics partially address this limitation, though they introduce hyperparameters governing the appropriate variance target. Recent work notes that geometric interpolation methods incorporating both magnitude and variance preservation consistently outperform naive averaging when source models exhibit low angular similarity, though the benefits diminish for models sharing high linear mode connectivity (Singh and Jaggi, 2020) where Euclidean assumptions remain approximately valid. While these geometric considerations inform how interpolation should proceed, an orthogonal question concerns what exactly should be interpolated, a perspective that motivates reconceptualizing fine-tuning through the lens of task vectors. Table 3 compares the weight-space averaging and geometric interpolation methods discussed above.
Table 3: Comparison of weight-space averaging and geometric interpolation methods for model merging.
@@
Element-wise arithmetic mean: $\theta_{\text{merged}}=\frac{1}{N}\sum_{i=1}^{N}\theta_{i}$ Simple; no hyperparameters; treats weight space as Euclidean Assumes equal parameter importance; suboptimal when models diverge significantly 2022
@@
Task Arithmetic
(Ilharco et al., 2023)
Compute and combine task vectors $\tau=\theta_{\text{ft}}-\theta_{\text{pre}}$ Modular; supports task addition, negation, and composition Task interference; requires shared pretrained initialization 2023
TIES-Merging
(Yadav et al., 2023)
Trim redundant params, elect signs, merge disjoint sets Reduces parameter interference; resolves sign conflicts Hyperparameter sensitive (trim ratio); additional computation 2023
Latent Space Merging
(Li et al., 2023a)
Interpolation in activation/latent space rather than weights Bypasses weight space geometry issues; often outperforms weight averaging May require data access; increased inference complexity 2024
@@
The methods examined in this section demonstrate that weight-space averaging, whether through uniform linear combinations, Fisher-weighted importance schemes, trajectory-based accumulation along optimization paths, or geometry-aware interpolation on parameter manifolds, provides a powerful framework for combining multiple models without additional training, with performance gains stemming from the exploitation of loss landscape connectivity and mode connectivity properties. While these approaches focus on holistically merging entire model checkpoints through various averaging strategies, the next section examines task vector arithmetic and sparsification-enhanced methods, which shift the perspective from combining full parameter sets to treating fine-tuning-induced weight differences as modular, composable units supporting addition, negation, and scaling for targeted multi-task composition.
@@ 4 Task Vector Arithmetic and Sparsification-Enhanced Methods
We now turn to a different paradigm: isolating and manipulating the task-specific knowledge encoded during fine-tuning. Task vector arithmetic, introduced by Ilharco et al. (2023), reconceptualizes model merging by treating the weight differences between fine-tuned models and their shared pre-trained ancestor as composable vectors that support algebraic operations such as addition, negation, and scaling. We also cover sparsification-enhanced variants such as TIES-Merging (Yadav et al., 2023) and DARE (Yu et al., 2023) that address parameter interference and sign conflicts.
@@ 4.1 Task Vector Formulation and Conceptual Framework
The task vector framework formalizes how fine-tuning encodes task-specific knowledge within neural network parameters. Ilharco et al. (2023) introduced this foundational framework by proposing that the knowledge acquired during fine-tuning can be isolated, manipulated, and transferred through simple algebraic operations on model weights. This formulation enables model specialization becomes amenable to systematic composition and modification.
Formally, given a pretrained model with parameters $\theta_{\text{pre}}$ and a model fine-tuned on task $t$ with resulting parameters $\theta_{t}$, the task vector $\tau_{t}$ is defined as the element-wise difference:
$\tau_{t}=\theta_{t}-\theta_{\text{pre}}$ (3)
This definition captures the parametric displacement induced by fine-tuning, effectively encoding the directional shift in weight space that transforms a general-purpose foundation model into a task-specialized variant. Based on this formulation, Ilharco et al. (2023) defined three fundamental arithmetic operations that constitute a versatile toolkit for training-free model editing:
@@
$\theta_{\text{multi}}=\theta_{\text{pre}}+\sum_{i=1}^{n}\lambda_{i}\tau_{i}$ (4)
where $\lambda_{i}$ denotes the scaling coefficient for task $i$. This operation enables the creation of multi-task models without joint training, assuming the task vectors reside in approximately orthogonal subspaces.
@@
$\theta_{\text{forget}}=\theta_{\text{pre}}-\lambda\tau$ (5)
@@
$\theta_{\text{scaled}}=\theta_{\text{pre}}+\alpha\tau$ (6)
@@
Task Analogy (Relationship Transfer). Extending the vector algebra metaphor, task analogy enables transferring the relationship between two tasks to a third domain. Given task vectors $\tau_{A}$, $\tau_{B}$, and a target task vector $\tau_{C}$, the analogy operation applies the directional difference between two tasks to a third:
$\theta_{\text{analogy}}=\theta_{\text{pre}}+\tau_{C}+(\tau_{B}-\tau_{A})$ (7)
@@
This operation mirrors classical word analogy reasoning (e.g., “king - man ++ woman == queen”) in embedding spaces, but operates directly on model parameters. As illustrated in Figure 3(d), task analogy enables compositional transfer of inter-task relationships without additional training.
@@
The conceptual elegance of this formulation lies in its interpretation: fine-tuning is reconceptualized not as an opaque optimization process but as the application of a structured transformation that can be decomposed, analyzed, and recombined with other such transformations.
@@
The task vector paradigm derives its power from treating these parameter deltas as first-class mathematical objects inhabiting a vector space (Ilharco et al., 2023). Under this interpretation, the weight space of neural networks admits linear algebraic operations that preserve semantic meaning. When a model is fine-tuned on a specific capability, whether instruction following, mathematical reasoning, or domain expertise, the resulting task vector encodes a compact representation of that capability that remains meaningful when extracted from its original context and applied to compatible base models.
@@
This formulation maintains deep connections to the theoretical foundations of mode connectivity discussed in earlier sections. The success of task vector arithmetic implicitly relies on the phenomenon of linear mode connectivity (Singh and Jaggi, 2020) between fine-tuned models sharing a common pretrained initialization (Garipov et al., 2018). When fine-tuned models reside within the same loss basin as their shared pretrained ancestor, the linear interpolation paths between them traverse regions of relatively low loss, enabling meaningful combination through arithmetic operations on their corresponding task vectors.
@@
The requirements for task vector compatibility impose important constraints on practical applications. As emphasized in (Li et al., 2023a), task vector operations are inherently limited to models with identical architectures since they rely on element-wise parameter correspondence. Both source models must derive from the same pretrained checkpoint, ensuring that their parameter spaces are aligned and that the extracted task vectors encode semantically coherent transformations. This architectural constraint reflects the fundamental observation that task vectors represent relative displacements within a shared reference frame established by the common initialization.
@@
The magnitude and distribution of task vector components carry meaningful information about the nature of fine-tuning adaptations. Ilharco et al. (2023) investigated how different fine-tuning procedures produce task vectors with distinct statistical properties, affecting downstream merging behavior. Empirical analyses reveal that task vectors exhibit considerable sparsity, with relatively few parameters undergoing substantial modification during fine-tuning while the majority remain close to their pretrained values (Yu et al., 2023). This sparsity observation has motivated various enhancement strategies, though the immediate practical power of task vectors emerges from the fundamental algebraic operations they support, addition, negation, and scaling, which together constitute a versatile toolkit for training-free model editing. Similar to how knowledge distillation (Kim et al., 2023; Hinton et al., 2015) transfers knowledge between models through output matching, task vectors enable knowledge transfer through direct parameter manipulation. Figure 3 illustrates the three primary task vector operations.
@@ 4.2 Basic Arithmetic Operations for Model Editing
The three primary operations—addition, negation, and scaling—transform task vectors into versatile tools for training-free model editing.
@@
Task Vector Addition represents the most widely employed operation, enabling the combination of multiple specialized capabilities into a unified model. Given a set of $n$ task vectors $\{\tau_{1},\tau_{2},\ldots,\tau_{n}\}$ extracted from models fine-tuned on distinct tasks, the merged model parameters are computed as $\theta_{\text{merged}}=\theta_{\text{pre}}+\sum_{i=1}^{n}\lambda_{i}\tau_{i}$, where $\lambda_{i}$ denotes the scaling coefficient for task $i$. The seminal work Ilharco et al. (2023) demonstrated that this simple additive combination successfully transfers capabilities across diverse task categories, including instruction following, mathematical reasoning, and domain-specific knowledge (Xiao et al., 2023) adaptation in LLMs. The effectiveness of addition relies critically on the assumption that task-specific knowledge encoded in different task vectors occupies approximately orthogonal subspaces of the parameter manifold, allowing superposition without destructive interference. (Li et al., 2023a) provides comprehensive empirical evidence that task addition achieves competitive performance across numerous NLP benchmarks when source tasks exhibit complementary rather than conflicting requirements.
@@
Task Vector Negation offers an effective capability for targeted behavior removal without requiring access to negative training examples or explicit unlearning procedures. By computing $\theta_{\text{edited}}=\theta_{\text{pre}}-\lambda\tau_{\text{unwanted}}$, practitioners can attenuate or eliminate specific model behaviors such as toxic language generation, demographic biases, or domain-specific knowledge (Xiao et al., 2023) that should be restricted. Ilharco et al. (2023) originally demonstrated this capability by showing that negating task vectors extracted from models fine-tuned on toxic content datasets reduced harmful generation rates while preserving general linguistic competence. The mathematical intuition underlying negation follows from the interpretation of task vectors as directional displacements in capability space; reversing the direction of displacement moves the model away from the associated capability manifold. However, Ilharco et al. (2023) revealed important nuances regarding fairness implications, demonstrating that naive negation operations can inadvertently introduce new biases or disproportionately affect model behavior across different demographic groups. This finding underscores the need for careful analysis when employing negation for bias mitigation purposes.
@@
Scaling Coefficients provide fine-grained control over the magnitude of task contributions, enabling balance competing objectives and prevent any single task from dominating the merged representation. The choice of scaling parameter $\lambda$ presents a fundamental tradeoff: larger values increase task-specific performance but risk catastrophic interference with the pretrained model’s general capabilities, while smaller values preserve base model competence at the cost of diminished task specialization. Unlike Ilharco et al. (2023), which primarily explored uniform scaling across tasks, subsequent investigations have demonstrated the benefits of task-specific coefficient tuning. Tang et al. (2024c) showed that operating in the tangent space of neural network weights enables more principled coefficient selection by linearizing the relationship between scaling magnitude and performance outcomes.
@@
The analogy between task arithmetic and classical linear algebra operations provides both conceptual clarity and practical limitations. While vector addition in Euclidean space enjoys exact superposition properties, task vector addition operates in a highly nonlinear parameter space where interactions between components can produce emergent behaviors not predicted by simple linear models  (Li et al., 2023a). These nonlinear interactions manifest as parameter-level conflicts that can substantially degrade merged model performance, interference phenomena that demand systematic examination to enable principled mitigation strategies.

### 30 Mar 2026 v2

@@ 3.4 Geometric Interpolation and Manifold-Aware Methods
Since Euclidean assumptions become problematic for divergent models, geometric interpolation techniques offer structured alternatives that account for the curved structure of parameter spaces. When combining two parameter vectors $\theta_{A}$ and $\theta_{B}$ via linear interpolation $\theta_{\text{merged}}=(1-\alpha)\theta_{A}+\alpha\theta_{B}$, the resulting vector’s magnitude typically satisfies $\|\theta_{\text{merged}}\|<(1-\alpha)\|\theta_{A}\|+\alpha\|\theta_{B}\|$ unless the vectors are perfectly aligned, leading to systematic magnitude shrinkage that can degrade model representations. This phenomenon proves particularly acute when merging models with markedly different weight configurations.
Spherical Linear Interpolation (SLERP) tackles magnitude preservation by constraining the interpolation path to lie on the hypersphere connecting two points in weight space. For two parameter vectors $\theta_{A}$ and $\theta_{B}$, SLERP computes the interpolated point as:
@@
where $\Omega=\arccos(\theta_{A}\cdot\theta_{B}/\|\theta_{A}\|\|\theta_{B}\|)$ represents the angular distance between the two vectors. This formulation ensures that the interpolated vector maintains constant magnitude throughout the interpolation trajectory, traversing a geodesic arc rather than cutting through the interior of the sphere. Empirical evidence has shown that SLERP-based merging preserves representational geometry more faithfully than linear averaging, particularly for attention weight matrices where directional information carries semantic significance.
@@
In practice, applying SLERP to neural network merging requires careful consideration of granularity. Layer-wise application, where SLERP operates independently on each layer’s parameter tensor, has emerged as the dominant approach, as global application across flattened parameter vectors fails to capture the hierarchical structure of deep networks. Layer-wise SLERP yields superior capability retention compared to uniform linear averaging when merging models fine-tuned on semantically distant tasks, attributed to the preservation of layer-specific activation statistics.
@@
Beyond spherical interpolation, manifold-aware methods encompass techniques that explicitly model the non-Euclidean structure of parameter spaces. Latent space interpolation approaches project weight configurations into lower-dimensional manifolds where interpolation can be performed with reduced interference before mapping back to the original parameter space. This approach acknowledges that the effective dimensionality of viable weight configurations is far lower than the ambient parameter count, allowing interpolation paths that remain within the manifold of high-performing solutions.
@@
Variance preservation presents additional challenges beyond magnitude shrinkage. When averaging independent parameter vectors, the variance of the resulting distribution decreases proportionally to the number of models merged, potentially collapsing learned features toward uninformative mean values. Rescaling strategies that adjust merged parameters to match the original variance statistics partially mitigate this limitation, though they introduce hyperparameters governing the appropriate variance target. Geometric interpolation methods incorporating both magnitude and variance preservation consistently outperform naive averaging when source models exhibit low angular similarity, though the benefits diminish for models sharing high linear mode connectivity (Frankle et al., 2020) where Euclidean assumptions remain approximately valid. Yet all methods discussed so far treat model parameters as quantities to be combined, without explicitly modeling what changed during fine-tuning. The next section introduces a complementary perspective, namely reconceptualizing fine-tuning as vector displacement in weight space, which enables targeted operations such as capability addition, negation, and interference-aware composition. Table 3 compares the weight-space averaging and geometric interpolation methods discussed above.
Table 3: Weight-space averaging and geometric interpolation methods for model merging. Methods range from simple uniform averaging to manifold-aware interpolation, progressively incorporating parameter importance (Fisher), temporal structure (SWA), and geometric properties (SLERP, manifold) to improve merge quality. “Key Approach” summarizes the merging mechanism; “Limitations” notes the primary constraint for each method.
@@
Element-wise arithmetic mean: $\theta_{\text{merged}}=\frac{1}{N}\sum_{i=1}^{N}\theta_{i}$ Simple; no hyperparameters; treats weight space as Euclidean Assumes equal parameter importance; suboptimal when models diverge substantially 2022
@@
Latent Space Merging Interpolation in activation/latent space rather than weights Bypasses weight space geometry issues; often outperforms weight averaging May require data access; increased inference complexity 2024
@@
As reviewed above, weight-space averaging, whether through uniform linear combinations, Fisher-weighted importance schemes, trajectory-based accumulation along optimization paths, or geometry-aware interpolation on parameter manifolds, provides a powerful basis for combining multiple models without additional training, with performance gains stemming from the exploitation of loss surface connectivity and mode connectivity properties. Recent advances continue to refine these foundations along several complementary axes, revealing a consistent trend, namely that the field is moving from uniform, method-agnostic averaging toward context-sensitive strategies that adapt merging behavior to the specific characteristics of the models and layers being combined.
A first line of work pursues data-free adaptive merging that replaces uniform coefficients with automatically derived, layer-wise or parameter-wise weights. FroM (Li et al., 2025e) uses Frobenius norm statistics, MAGIC (Li et al., 2025c) exploits magnitude calibration, and Orthogonal Model Merging (Yang et al., 2026c) preserves orthogonality during combination. The shared insight across these methods is that different layers contribute unequally to task-specific behavior, and norm-based or structural heuristics can approximate this heterogeneity without requiring any training data.
A second direction extends averaging from Euclidean weight space to richer geometric structures. ACE-Merging (Xu et al., 2026) uses adaptive covariance estimation, while Functionality-Oriented Merging (Wang et al., 2026b) operates on the Fisher–Rao manifold. These approaches trade computational simplicity for tighter alignment with the functional geometry of neural networks, yielding improvements particularly when source models have undergone substantial fine-tuning that violates local linearity assumptions.
Third, importance-weighted refinements such as Dynamic Fisher-weighted Merging (Lee et al., 2025b), which optimizes Fisher weights via Bayesian optimization, and RegMean++ (Nguyen et al., 2025c) revisit the classical idea of curvature-based weighting with modern optimization tools, reducing the sensitivity of Fisher-based methods to hyperparameter choices.
Fourth, a growing body of work integrates merging into the training pipeline itself rather than applying it only post hoc. WSM (Tian et al., 2025) introduces decay-free learning rate schedules via checkpoint merging, Anytime Pretraining (Meterez et al., 2026) develops horizon-free schedules with weight averaging, and Mashup Learning (Vaina et al., 2026) accelerates fine-tuning by remixing checkpoints. This trend blurs the boundary between optimization and merging, suggesting that future training protocols may be co-designed with merging objectives.
Finally, scalable and training-free approaches target practical deployment at LLM scale. Training-free LLM Merging (Fu et al., 2025c) proposes hierarchical pruning and scaling across model and layer levels for multi-task merging on LLaMA-3 and Qwen, while Harmonizing Diverse Models (Peng et al., 2025) introduces consistency-aware layer-wise weights computed from intermediate activations to improve generation coherence in RAG systems. Further works tackle weight averaging from the perspective of pretraining data mixture optimization (Wang et al., 2026a; Li et al., 2026b; Zhu et al., 2026; Tao et al., 2025), parameter-efficient checkpoint merging (Yu and Choi, 2025), and bagging-based merging for robust text embeddings (Zhang et al., 2026). While these approaches focus on holistically merging entire model checkpoints through various averaging strategies, the next section examines task vector arithmetic and sparsification-enhanced methods, which shift the perspective from combining full parameter sets to treating fine-tuning-induced weight differences as modular, composable units supporting addition, negation, and scaling for targeted multi-task composition.
@@ 4 Task Vector Arithmetic and Sparsification-Enhanced Methods
Rather than treating model parameters as interchangeable quantities to be averaged, an alternative perspective focuses on isolating and manipulating the task-specific knowledge encoded during fine-tuning. Task vector arithmetic, introduced by Ilharco et al. (2023), reconceptualizes model merging by treating the weight differences between fine-tuned models and their shared pretrained ancestor as composable vectors that support algebraic operations such as addition, negation, and scaling. This section also covers sparsification-enhanced variants such as TIES-Merging (Yadav et al., 2023) and DARE (Yu et al., 2024) that tackle parameter interference and sign conflicts.
@@ 4.1 Task Vector Formulation and Conceptual Framework
The task vector framework formalizes how fine-tuning encodes task-specific knowledge within neural network parameters. Ilharco et al. (2023) introduced this foundational framework by proposing that the knowledge acquired during fine-tuning can be isolated, manipulated, and transferred through simple algebraic operations on model weights, making model specialization amenable to systematic composition and modification.
Formally, given a pretrained model with parameters $\theta_{\text{pre}}$ (denoted $\theta_{0}$ in Section 2) and a model fine-tuned on task $t$ with resulting parameters $\theta_{t}$, the task vector $\tau_{t}$ is defined as the element-wise difference:
$\tau_{t}=\theta_{t}-\theta_{\text{pre}}.$ (3)
This definition captures the parametric displacement induced by fine-tuning, effectively encoding the directional shift in weight space that transforms a general-purpose foundation model into a task-specialized variant. Based on this formulation, Ilharco et al. (2023) defined three fundamental arithmetic operations that constitute a versatile toolkit for training-free model editing, namely
@@
$\theta_{\text{multi}}=\theta_{\text{pre}}+\sum_{i=1}^{n}\lambda_{i}\tau_{i},$ (4)
where $\lambda_{i}$ denotes the scaling coefficient for task ii. This operation supports the creation of multi-task models without joint training, assuming the task vectors reside in approximately orthogonal subspaces.
@@
$\theta_{\text{forget}}=\theta_{\text{pre}}-\lambda\tau,$ (5)
@@
$\theta_{\text{scaled}}=\theta_{\text{pre}}+\alpha\tau,$ (6)
@@
Task Analogy (Relationship Transfer). Extending the vector algebra metaphor, task analogy enables transferring the relationship between two tasks to a third domain. Given task vectors $\tau_{A}$, $\tau_{B}$, and a target task vector $\tau_{C}$, the analogy operation applies the directional difference between two tasks to a third as follows.
$\theta_{\text{analogy}}=\theta_{\text{pre}}+\tau_{C}+(\tau_{B}-\tau_{A}).$ (7)
@@
This operation mirrors classical word analogy reasoning (e.g., “king - man ++ woman == queen”) in embedding spaces, but operates directly on model parameters. As illustrated in Figure 3(d), task analogy permits compositional transfer of inter-task relationships without additional training.
@@
Conceptually, this formulation’s elegance lies in its interpretation. Fine-tuning is reconceptualized not as an opaque optimization process but as the application of a structured transformation that can be decomposed, analyzed, and recombined with other such transformations.
@@
The task vector framework derives its power from treating these parameter deltas as first-class mathematical objects inhabiting a vector space (Ilharco et al., 2023). Under this interpretation, the weight space of neural networks admits linear algebraic operations that preserve semantic meaning. When a model is fine-tuned on a specific capability, whether instruction following, mathematical reasoning, or domain expertise, the resulting task vector encodes a compact representation of that capability that remains meaningful when extracted from its original context and applied to compatible base models.
@@
This formulation maintains deep connections to the theoretical foundations of mode connectivity discussed in earlier sections. The success of task vector arithmetic implicitly relies on the phenomenon of linear mode connectivity (Frankle et al., 2020) between fine-tuned models sharing a common pretrained initialization. When fine-tuned models reside within the same loss basin as their shared pretrained ancestor, the linear interpolation paths between them traverse regions of relatively low loss, permitting meaningful combination through arithmetic operations on their corresponding task vectors.
@@
Task vector compatibility imposes important constraints on practical applications. Task vector operations are inherently limited to models with identical architectures since they rely on element-wise parameter correspondence. Both source models must derive from the same pretrained checkpoint, ensuring that their parameter spaces are aligned and that the extracted task vectors encode semantically coherent transformations. This architectural constraint reflects the key observation that task vectors represent relative displacements within a shared reference frame established by the common initialization.
@@
Both the magnitude and distribution of task vector components carry meaningful information about the nature of fine-tuning adaptations. Ilharco et al. (2023) investigated how different fine-tuning procedures produce task vectors with distinct statistical properties, affecting downstream merging behavior. Empirical analyses reveal that task vectors exhibit considerable sparsity, with relatively few parameters undergoing large modifications during fine-tuning while the majority remain close to their pretrained values (Yu et al., 2024). This sparsity observation has motivated various enhancement strategies, though the immediate practical power of task vectors emerges from the fundamental algebraic operations they support, addition, negation, and scaling, which enable training-free model editing without access to training data. Similar to how knowledge distillation (Hinton et al., 2015) transfers knowledge between models through output matching, task vectors enable knowledge transfer through direct parameter manipulation. Figure 3 illustrates the three primary task vector operations.
@@ 4.2 Basic Arithmetic Operations for Model Editing
The three primary operations, addition, negation, and scaling, transform task vectors into versatile tools for training-free model editing.
@@
Task Vector Addition represents the most widely used operation, enabling the combination of multiple specialized capabilities into a unified model. Given task vectors $\{\tau_{1},\tau_{2},\ldots,\tau_{n}\}$ from $n$ fine-tuned models, the merged parameters follow Eq. 4. The seminal work Ilharco et al. (2023) showed that this simple additive combination successfully transfers capabilities across diverse task categories, including instruction following, mathematical reasoning, and domain-specific knowledge adaptation in LLMs. The effectiveness of addition relies critically on the assumption that task-specific knowledge encoded in different task vectors occupies approximately orthogonal subspaces of the parameter manifold, allowing superposition without destructive interference. Empirical evidence indicates that task addition achieves competitive performance across numerous NLP benchmarks when source tasks exhibit complementary rather than conflicting requirements.
@@
Task Vector Negation offers an effective mechanism for targeted behavior removal without requiring access to negative training examples or explicit unlearning procedures. By computing $\theta_{\text{edited}}=\theta_{\text{pre}}-\lambda\tau_{\text{unwanted}}$, practitioners can attenuate or eliminate specific model behaviors such as toxic language generation, demographic biases, or restricted domain knowledge (Ilharco et al., 2023). The mathematical intuition follows from interpreting task vectors as directional displacements in capability space; reversing the direction moves the model away from the associated capability manifold. However, care must be taken, as the interaction between negated and retained capabilities may produce unintended side effects. Over-aggressive negation (large $\lambda$) can degrade general capabilities beyond the targeted behavior, and the impact may vary across demographic groups (Ilharco et al., 2023).
Several extensions have refined the negation operation for specific applications. Elastic Weight Removal (Daheim et al., 2024) uses Fisher information to weigh parameter importance when subtracting a “negative expert” fine-tuned on hallucinated outputs, enabling more precise removal of unfaithful generation while preserving abstractive capabilities. Separate the Wheat from the Chaff (Hu et al., 2024) improves negation through an extraction-before-subtraction approach that isolates deficiency-specific parameters before removal. Controlled text generation via language model arithmetic (Dekoninck et al., 2024) extends the arithmetic metaphor to the inference level, composing multiple LLMs through algebraic operations on output distributions rather than model parameters for fine-grained control over generation attributes. Forgetting before Learning (Ni et al., 2024) demonstrates that parametric arithmetic can first subtract parameters encoding obsolete information before fine-tuning on new knowledge.
@@
Scaling Coefficients provide fine-grained control over the magnitude of task contributions, enabling practitioners to balance competing objectives and prevent any single task from dominating the merged representation. The choice of scaling parameter $\lambda$ presents a fundamental tradeoff, as larger values increase task-specific performance but risk catastrophic interference with the pretrained model’s general capabilities, while smaller values preserve base model competence at the cost of diminished task specialization. Unlike Ilharco et al. (2023), which primarily explored uniform scaling across tasks, subsequent investigations have revealed the benefits of task-specific coefficient tuning. Tang et al. (2024d) showed that partially linearizing adapter modules and applying task arithmetic over the linearized adapters enables more principled coefficient selection by improving the linearity assumptions underlying weight-space composition.
@@
An analogy between task arithmetic and classical linear algebra operations provides both conceptual clarity and practical limitations. While vector addition in Euclidean space enjoys exact superposition properties, task vector addition operates in a highly nonlinear parameter space where interactions between components can produce emergent behaviors not predicted by simple linear models. These nonlinear interactions manifest as parameter-level conflicts that can substantially degrade merged model performance, interference phenomena that demand systematic examination to support principled mitigation strategies.